# Token Embeddings Violate the Manifold Hypothesis

### v2

@@ A.7. Tables of manifold rejections in all four LLMs
7 porque 0.0000128 Must start a word
8 Leopold 0.0000135 Must start a word
9 Rem 0.0000232 Must start a word
10 Ern 0.0000265 Must start a word
11 though 0.0000540 Must start a word
12 č 0.0000619 Must start a word
13 endl 0.0000683 Must start a word
14 corrected 0.0000738 Must start a word
15 tw 0.0000788 Must start a word
16 ères 0.000115
17 grep 0.000151
18 kind 0.000202 Must start a word
19 pax 0.000243
20 Tri 0.000250 Must start a word
21 spielte 0.000258 Must start a word
22 Heaven 0.000262 Must start a word
23 d 0.000289
24 ¡Cyrillic¿ 0.000320
25 Card 0.000335 Must start a word
26 Wall 0.000351 Must start a word
27 periodic 0.000359 Must start a word
28 ucht 0.000463
29 jeu 0.000553 Must start a word
30 slightly 0.000807 Must start a word
31 passion 0.000823 Must start a word
32 Charles 0.000833 Must start a word
33 victory 0.000906 Must start a word
Table 7. Rejections of the manifold hypothesis for Mistral7B
Number Token $p$-value Comment
1 CV 0.000000241 Must start a word
2 ember 0.00000199
3 ¡Korean character¿ 0.00000224
4 ĭ 0.00000770
5 rise 0.0000283 Must start a word
6 poured 0.0000305 Must start a word
7 aimed 0.0000364 Must start a word
8 allocate 0.0000553 Must start a word
9 UINT 0.0000684
10 monitor 0.0000767 Must start a word
11 ¡Cyrillic¿ 0.000105
[…]
13 ¡Chinese character¿ 0.000133
14 Lat 0.000139 Must start a word
15 Vel 0.000148
16 challenges 0.000178 Must start a word
17 ¡Chinese character¿ 0.000200
18 Od 0.000222 Must start a word
19 neys 0.000244
20 selon 0.000244 Must start a word
21 measures 0.000268 Must start a word
22 assumes 0.000321 Must start a word
23 Et 0.000338 Must start a word
24 Rot 0.000438
25 ¡Chinese character¿ 0.000538
26 olutely 0.000540
27 HO 0.000567
28 Solutions 0.000598 Must start a word
29 * 0.000658
30 Condition 0.000665
31 zt 0.000677
32 iced 0.000696
33 URL 0.000722
34 SEO 0.000728 Must start a word
35 weigh 0.000728 Must start a word
36 änge 0.000765
37 Vic 0.000777 Must start a word
38 apparent 0.000780 Must start a word
39 ¡Chinese character¿ 0.000896
40 ¡0xE3¿ 0.000921
Table 8. Rejections of the manifold hypothesis for Pythia6.9B (1 of 2)
Number Token $p$-value Comment
1 psychotic 0.000000161 Must start a word
2 \\]( 0.00000147
3 1990 0.00000231
4 tiene 0.00000282 Must start a word
5 ucky 0.0000134
6 Hebrew 0.0000141 Must start a word
7 mess 0.0000184
8 reasonable 0.0000300 Must start a word
9 collector 0.0000311 Must start a word
10 embodiments 0.0000428 Must start a word
11 Inhibition 0.0000593 Must start a word
12 Hers 0.0000691 Must start a word
13 odge 0.000120
14 áĢIJ 0.000144
15 ÏĢο 0.000165
16 198 0.000170 Must start a word
17 scare 0.000192 Must start a word
18 ]{}\\_[ 0.000202
19 LIMITED 0.000212 Must start a word
20 hop 0.000232 Must start a word
21 BIO 0.000246
22 raison 0.000252 Must start a word
23 inside 0.000254
24 st 0.000261
25 waving 0.000286 Must start a word
26 Austria 0.000287 Must start a word
27 ny 0.000313
Table 9. Rejections of the manifold hypothesis for Pythia6.9B (2 of 2)
Number Token $p$-value Comment
28 bags 0.000326
29 medium 0.000372
30 incent 0.000378 Must start a word
31 carboh 0.000407 Must start a word
32 secret 0.000451 Must start a word
33 ¡Unicode characters¿ 0.000468
34 ÏĢε 0.000519 Must start a word
35 120 0.000536
36 wastes 0.000565 Must start a word
37 ressor 0.000570
38 Lip 0.000597 Must start a word
39 preview 0.000650
40 redund 0.000653 Must start a word
41 ridic 0.000673 Must start a word
42 Ramirez 0.000698 Must start a word
43 omatic 0.000707
44 Gloria 0.000733 Must start a word
45 eton 0.000780
46 chairman 0.000789 Must start a word
47 Supreme 0.000859 Must start a word
48 grazing 0.000890 Must start a word
49 activation 0.000892 Must start a word
50 imming 0.000914
51 radiological 0.000931 Must start a word
52 branc 0.000941 Must start a word
53 \$\\—\\ 0.000942 Must start a word
54 DEFAULT 0.000992

### v3

@@ A.9 Tables of manifold rejections in all four LLMs
13 31517 0.000133 Chinese
14 7402 Lat 0.000139 Must start a word
15 24351 Vel 0.000148
16 10369 challenges 0.000178 Must start a word
17 30085 0.000200 Chinese
18 11424 Od 0.000222 Must start a word
19 26983 neys 0.000244
20 23851 selon 0.000244 Must start a word
21 10582 measures 0.000268 Must start a word
22 27460 assumes 0.000321 Must start a word
23 13542 Et 0.000338 Must start a word
24 13722 Rot 0.000438
25 31097 0.000538 Chinese
26 25244 olutely 0.000540
27 4104 HO 0.000567
28 24806 Solutions 0.000598 Must start a word
29 29587 * 0.000658
30 12862 Condition 0.000665
31 3903 zt 0.000677
32 3903 iced 0.000696
33 5343 URL 0.000722
34 21518 SEO 0.000728 Must start a word
35 19929 weigh 0.000728 Must start a word
36 19929 änge 0.000765
37 19929 Vic 0.000777 Must start a word
38 7566 apparent 0.000780 Must start a word
39 31780 0.000896 Chinese
40 230 <0xE3> 0.000921
Table 9: Rejections of the manifold hypothesis for Pythia6.9B (1 of 2)
Number Token ID Token $p$-value Comment
1 49503 psychotic 0.000000161 Must start a word
2 49503 \\]( 0.00000147
3 13435 1990 0.00000231
4 31918 tiene 0.00000282 Must start a word
5 12202 ucky 0.0000134
6 25260 Hebrew 0.0000141 Must start a word
7 46631 mess 0.0000184
8 5272 reasonable 0.0000300 Must start a word
9 25389 collector 0.0000311 Must start a word
10 24898 embodiments 0.0000428 Must start a word
11 47769 Inhibition 0.0000593 Must start a word
12 41264 Hers 0.0000691 Must start a word
13 13676 odge 0.000120
14 49669 0.000144 Turkish
15 16287 0.000165 Turkish
16 26397 198 0.000170 Must start a word
17 30272 scare 0.000192 Must start a word
18 41305 ]{}\\_[ 0.000202
19 25666 LIMITED 0.000212 Must start a word
20 5184 hop 0.000232 Must start a word
21 39253 BIO 0.000246
22 48089 raison 0.000252 Must start a word
23 40084 inside 0.000254
24 296 st 0.000261
25 34475 waving 0.000286 Must start a word
26 21753 Austria 0.000287 Must start a word
27 5134 ny 0.000313
Table 10: Rejections of the manifold hypothesis for Pythia6.9B (2 of 2)
Number Token ID Token $p$-value Comment
28 46728 bags 0.000326
29 28672 medium 0.000372
30 15210 incent 0.000378 Must start a word
31 21742 carboh 0.000407 Must start a word
32 4279 secret 0.000451 Must start a word
33 50231 0.000468 misc. Unicode
34 40146 0.000519 Turkish, Must start a word
35 8193 120 0.000536
36 47334 wastes 0.000565 Must start a word
37 32232 ressor 0.000570
38 22609 Lip 0.000597 Must start a word
39 42150 preview 0.000650
40 19886 redund 0.000653 Must start a word
41 41962 ridic 0.000673 Must start a word
42 48444 Ramirez 0.000698 Must start a word
43 8977 omatic 0.000707
44 46972 Gloria 0.000733 Must start a word
45 18807 eton 0.000780
46 16782 chairman 0.000789 Must start a word
47 6413 Supreme 0.000859 Must start a word
48 42351 grazing 0.000890 Must start a word
49 5743 activation 0.000892 Must start a word
50 34282 imming 0.000914
51 46513 radiological 0.000931 Must start a word
52 11363 branc 0.000941 Must start a word
53 47183 $\|\\$ 0.000942 Must start a word
54 25409 DEFAULT 0.000992
@@
A.10 Limitations of theory
Because our approach requires the distances between all pairs of tokens, to use the hypothesis tests presented in this paper, one needs to have access to the token embedding vectors. This is why our paper considers four open source models of moderate size – GPT2 (Radford et al. [2019]), Llemma7B (Azerbayev et al. [2024]), Mistral7B (Jiang et al. [2023]), and Pythia6.9B (Biderman et al. [2023]). While Robinson et al. [2025] does provide an approach to extract the token embeddings using systematic prompting, which would allow the application of our tests to proprietary models, their method requires exquisite control of the prompting and is computationally expensive.
The reach itself is difficult to estimate accurately from sampled data (Berenfeld et al. [2022], Rawson and Robinson [2021], Aamari and Levrard [2019]). Our implementation tests the radius below an estimated reach, determined by an inspection of the radius-vs-volume plots of a random sample of tokens. If the true reach is smaller than our estimates, the curvature of the manifolds that fit the data well, found by any method, will be high.